A guitar string is 90 cm long and has a mass of 3.6 g . The distance from the bridge to the support post is L=62cm, and the string is under a tension of 540 N .

What are the frequencies of the fundamental and first two overtones?

Express your answers using two significant figures. Enter your answers in ascending order separated by commas.
Answer #1

speed = sqrt(T/u)

where T = tension

u = mass per unit

v = sqrt(540 / (0.0036/0.9)

= 367.42 m/s

wavelength = 2L

= 2*0.62 = 1.24 m

speed = frequency * wavelength

v = f*\lambda

f = 367.42 / 1.24 = 296.3 Hz = 3.0 *10^2 Hz